OpenStudy (anonymous):

combine the following rational expressions. reduce all answers to lowest terms. (3)/(x+3) + (x)/(x+3)

OpenStudy (anonymous):

the denominators are the same, so add in the numerator \[\frac{3}{x+3}+\frac{x}{x+3}=\frac{x+3}{x+3}\]

OpenStudy (anonymous):

and since numerator and denominator are the same, you get 1